Circular No. 8700

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2576 OPHIUCHI = NOVA OPHIUCHI 2006 No. 2 P. Williams, Heathcote, New South Wales, reports his visual discovery

of a possible nova (mag 10.5) on Apr. 6.565 UT, noting it to be situated 4′.8 southeast of the nearby R CrB-type variable V517 Oph (itself invisible near minimum); Williams gave the position for the new star as α = 17^{h}15^{m}30^{s}, δ = −29^{o}09′30′′ (equinox 2000.0). E. O. Waagen, AAVSO, reports the following precise position for the new object from T. Krajci (Cloudcroft, NM): α = 17^{h}15^{m}33^{s}.00, δ = −29^{o}09′39′′.9 (20 UCAC2 stars used in the solution), adding that a nearby USNO-B1.0 star has position end figures 33^{s}.007, 36′′.84. Additional visual magnitude estimates: Apr. 6.590, 10.4 (Williams); 6.611, 10.4 (Williams); 7.154, 10.2 (P. Schmeer, Bischmisheim, Germany); 7.552, 9.2 (Williams); 7.575, 9.2 (Williams); 9.531, 9.3 (Williams).

H. Naito and S. Ozaki, Nishi-Harima Astronomical Observatory; and H. Yamaoka, Kyushu University, report that a low-resolution spectrum (resolution 1200 at Hα, range 420–680 nm) of this new object was obtained on Apr. 7.80 UT with the 2.0-m NAYUTA telescope under hazy conditions. The spectrum shows a flat continuum with prominent Balmer lines, showing a P-Cyg profile, and other lines in absorption, which suggests that it is indeed a classical nova caught at maximum light. The Hα peak shows a self-absorption feature; the FWHM of the Hα emission is 1470 km/s, and its absorption minimum is blueshifted by 1720 km/s from the emission peak obtained by Gaussian fitting. Yamaoka adds that M. Fujii (Ibara, Okayama, Japan) also took a spectrum (range 400–850 nm) of this object on Apr. 7.79, which shows the same features as noted above.

E. Kazarovets, Sternberg Astronomical Institute, Moscow University, reports that the permanent designation V2576 Oph has been assigned to this nova.

SUPERNOVA 2006bp IN NGC 3953 S. Nakano, Sumoto, Japan, reports the discovery by K. Itagaki (Teppo-

cho, Yamagata, Japan) of an apparent supernova (mag 16.7) on unfiltered 20-s CCD exposures taken around Apr. 9.60 UT (limiting mag 19.5) using a 0.60-m f/5.7 reflector. The new object has continued to brighten rapidly, as indicated by Itagaki’s additional magnitudes: Apr. 9.6136, 16.7; 9.6689, 16.4; 9.7383, 16.0; 9.7835, 15.8. SN 2006bp is located at α = 11^{h}53^{m}55^{s}.74, δ = +52^{o}21′09′′.4 (equinox 2000.0), which is 62′′ east and 93′′ north of the center of NGC 3953. Nothing was visible at this location on Itagaki’s images taken on 2005 Mar. 19 and 2006 Mar. 23 (limiting mag 19.0).

2006 April 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green